# Ngoc BuiY , Shubham SharmaJ, Simran LambaJ, Saumitra MishraJ, Rex YingY

### v1

@@ 1 Introduction
Modern large language models (LLMs) can, in principle, handle extremely long input contexts – some recent models support context windows of 128k tokens or more (yang2025qwen3; gao2024train). Yet, extending context length comes with steep computational costs. The self-attention mechanism has quadratic time complexity in sequence length, and storing the key-value (KV) cache for thousands of tokens can quickly exhaust GPU memory (wang2025llms; li2024survey). In practical deployments, the KV cache, which saves past key and value vectors to avoid re-computation, becomes a major memory and latency bottleneck for long-context inference. Decoupling resource usage from context length is therefore critical for enabling efficient and scalable applications such as long-horizon reasoning (chen2025towards) and lifelong agents (zheng2025lifelong; li2024larm).
@@
To address this challenge, recent work has explored memory-bounded LLMs that can operate effectively under constrained KV budgets (li2024survey). One line of research focuses on compression and quantization, aiming to reduce memory footprint by learning compact representations of past tokens rather than storing all keys and values explicitly (hooper2024kvquant; saxena2024eigen). These techniques are mostly effective during the prefill phase but scale poorly with generation length. Another line leverages attention sparsity to offload most of the cache to CPU or secondary storage, and retrieve only relevant segments on demand via similarity search (tang2024quest) or learned indices (gao2025seerattention). While offloading lowers the on-GPU footprint, it incurs nontrivial orchestration overhead that accumulates over long generations, undermining end-to-end throughput.
@@
A more common and direct approach to enforce a fixed memory budget is KV cache eviction, which directly drops certain tokens from the KV cache (xiao2023efficient). Many KV eviction strategies have been proposed to decide which tokens to remove. However, most of them are attention-guided heuristics: they track attention from new queries to cached tokens and retain those that are recently or frequently attended, adapting the cache to the current focus (zhang2023h2o; li2024snapkv; wang2025llms; liu2025can; ghadia2025dialogue; cai2025r). While being efficient, these methods assume that recent attention is a reliable proxy for future importance. This assumption often breaks for long-horizon generation and reasoning tasks: a token might be crucial much later, even if it has not been attended to in the recent past (jiang2024minference). Moreover, attention-based eviction can suffer from attention bias, e.g., the model might temporarily overlook a needed token due to a distracting context (shi2023large), causing it to be evicted prematurely. While some recent studies have attempted to learn better eviction decisions (chen2024nacl; zeng2024context), these methods typically scale poorly with sequence length and are therefore limited to the prefilling stage.
@@
In this work, we take a new perspective on the KV eviction problem. Rather than relying on the attention-guided importance, we propose to learn each token’s intrinsic importance at the time of its creation and use that as the basis for eviction. Intuitively, not all tokens are created equal: some carry significant semantic or task-related weight (e.g. a critical fact, a question being answered, or the first few “sink" tokens that often encode the topic or instructions), while others are relatively inconsequential (e.g. filler words, stopwords, or trivial arithmetic steps in a chain-of-thought). Moreover, the importance of tokens is not uniform across the network, but it varies systematically by layers and heads, reflecting their functional specializations (voita2019analyzing; wu2024retrieval).
@@
We posit that the contextual embedding of a token already encodes much of its long-term utility. We therefore introduce a retention gate that maps the token’s embedding and produces a scalar retention score $\beta\in[0,1]$ reflecting the token’s inherent importance for a specific layer and head. Especially, we design this retention score to decay exponentially as the context grows, mimicking the gradual forgetting of old information in human brains (ebbinghaus2013image). Thus, a highly important token will have $\beta\approx 1$ and retain a high score for a long time, whereas a token deemed unimportant will have $\beta$ closer to 0 and its influence will vanish quickly. We leverage this score to drive a simple eviction policy: whenever the number of cached tokens exceeds the budget $M$, we evict the token with the smallest current retention score. This approach, which we call Token RetentIon for Memory-bounded KV Cache (TRIM-KV), ensures that at all times, the cache is filled with the $M$ tokens judged most intrinsically important, with a preference toward more recently generated tokens.
@@
Results and Contributions. Through extensive experiments on long-context and long-generation benchmarks, we demonstrate that our learnable token retention approach substantially improves the performance of memory-bounded LLMs. On challenging mathematical reasoning datasets, GSM8K, MATH, AIME, a long procedural generation benchmark, LongProc, and a long-memory chat assistant benchmark, LongMemEval, our method consistently outperforms eviction baselines, even when those baselines use $4\times$ more KV budget, and deliver 58.4% pass@1 gain compared to the SOTA learnable KV retrieval baseline (gao2025seerattention), especially in low-memory regimes. Remarkably, in several settings, TRIM-KV even surpasses a full-cache model, suggesting that selective retention can function as an effective regularizer by suppressing noise from uninformative tokens.
We also present qualitative evidence that learned retention scores align with human intuition: the model tends to assign high scores to initial tokens and problem descriptions, and low scores to less meaningful punctuation. Notably, many behaviors reminiscent of common heuristics, such as keeping sink tokens, sliding windows, and gist compression (mu2023learning), emerge naturally and adaptively from our learned policy, without being hard-coded. Finally, we show that these learned retention scores can also act as a diagnostic tool for probing layer- and head-specific dynamics, providing a lightweight means to analyze and ultimately improve the interpretability of attention patterns.
@@ 2 Related Work
KV Cache Compression. As model sizes and context windows grow, optimizing KV-cache memory is increasingly critical. Prior work largely falls into three directions: (i) token eviction/merging (xiao2023efficient; li2024snapkv; zhang2023h2o; nawrot2024dynamic; zhang2024cam; qin2025cake; wang2025llms; liu2025can; park2025keydiff; cai2025r; park2025keydiff; kim2024infinipot), (ii) vector compression/quantization (hooper2024kvquant; liu2024kivi; yue2024wkvquant; sun2024shadowkv), and (iii) token retrieval (tang2024quest; liu2024retrievalattention; gao2025seerattention). While effective in many settings, vector compression and retrieval either discard fine-grained information or introduce nontrivial systems overhead (e.g., coordination and data movement) (li2024survey). Moreover, their memory and computation still scale with sequence length, making them inefficient for long-horizon generation applications. Token eviction offers a simple, memory-bounded alternative; however, most existing policies are heuristic and can significantly degrade performance, especially on long reasoning trajectories. Recent work has introduced learnable eviction policies (chen2024nacl; zeng2024context; huang2024locret), but these are primarily designed for the pre-filling stage and thus are not well suited to sustained long-horizon generation. We bridge this gap by introducing a learnable and efficient eviction policy designed for long-horizon LLM inference under fixed memory budgets.
Forgetting in Language Models. A key limitation of vanilla self-attention is the lack of an explicit forgetting mechanism, forcing the model to carry potentially irrelevant information and making long-context processing inefficient. Early work tackled this by replacing quadratic attention with linearized and recurrent variants (katharopoulos2020transformers; wang2020linformer; sun2023retentive; yang2023gated; yang2024parallelizing) that summarize the past into a fixed-size state, often a single vector. While computationally attractive, such heavy compression can degrade performance on tasks requiring long-range memory. Follow-up studies (behrouz2024titans; sun2024learning; karamitrellis; karami2025lattice) increase memory capacity by replacing this hidden vector with a more expressive neural state. Complementary lines of work retain softmax attention but enforce forgetting by modifying attention logits (lin2025forgetting) or imposing trainable sparsity patterns (yuan2025native). However, these approaches typically alter attention dynamics substantially and thus require training models from scratch. This incurs significant training cost and leaves their scalability to contemporary LLM sizes uncertain. In contrast, we introduce a plug-in forgetting mechanism for pretrained LLMs that converts them into memory-bounded models, providing long-context efficiency without retraining from scratch.
@@ 3.1 Transformers with Self-Attention
where $\mathbf{q},\mathbf{k},\mathbf{v}$ are query, key, and value states, respectively, and $\mathbf{W}_{Q},\mathbf{W}_{K},\mathbf{W}_{V}\in\mathbb{R}^{d\times d}$ are linear transformation weights. Here, we assume a single-head attention layer and omit the scaling factor $1/\sqrt{d}$ for simplicity. The sequence of key-value pairs $\{(\mathbf{k}_{i},\mathbf{v}_{i})\}_{i}$ is the in-context memory of the LLM. During the autoregressive decoding, we typically generate one token at a time and cache the running key-value pair $(\mathbf{k}_{t},\mathbf{v}_{t})$ to our in-context memory to avoid recomputation. However, this vanilla caching approach leads to a linear increase in memory footprint with the sequence length, while computation grows quadratically (keles2023computational). This reduces efficiency when handling long-context inputs and extended generation tasks.
@@ 3.2 Revisiting KV Cache Eviction
$\mathbf{o}_{t}^{\prime}=\sum_{i=1}^{t}\frac{\exp\left({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\alpha_{ti}}\mathbf{q}_{t}^{\top}\mathbf{k}_{i}\right)}{\sum_{j=1}^{t}\exp\left({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\alpha_{tj}}\mathbf{q}_{t}^{\top}\mathbf{k}_{j}\right)}\mathbf{v}_{i}\quad\text{where}\quad~\alpha_{ti}\in\{0,1\}~\text{and}~\alpha_{ti}\geq\alpha_{t+1,i},~~\forall i,t.\vskip-11.38109pt$ (1)
@@
Solving the above constrained optimization at every time step $t$ is impractical due to its combinatorial nature and efficiency requirements of LLM inference in real-world applications. Most existing approaches  (xiao2023efficient; han2023lm; zhang2023h2o; li2024snapkv; cai2025r; ghadia2025dialogue) opt to determine $\alpha$ heuristically while we focus on a learnable eviction method.
@@ 4.1 Selective In-Context Memory via Retention-Gated Attention
We introduce retention-gated attention, a trainable mechanism that mimics the information loss induced by inference-time eviction. From the formulation (1), the sequence $\alpha_{ii},\alpha_{(i+1)i},\ldots,\alpha_{ti}$ represents how token $i$ is retained in the attention computation over time. Retention begins at $1$ and then abruptly drops to 0 once the token is evicted. While this binary behavior matches the inference stage, it poses challenges for learning: the signal is discrete, non-differentiable, thus providing no gradients for optimization. To remedy this, we replace the hard binary variable $\alpha$ with a smooth, monotonically decreasing function that models the gradual decay of importance while enabling gradient-based training. A natural candidate is the sigmoid function, $\mathrm{bar}{\alpha}_{ti}=1/(1+\exp(f(\mathbf{x}_{i},t)))$, which predicts the time at which the token is evicted. However, this design suffers from two drawbacks: (i) the domain of $f$ is unnormalized since the sequence length is unknown during decoding, and (ii) the sigmoid flattens across most of its range, producing negligible variation between steps and leading to vanishing gradients during training.
@@
$\mathbf{q}_{t}=\mathbf{W}_{Q}\mathbf{x}_{t},\mathbf{k}_{t}=\mathbf{W}_{K}\mathbf{x}_{t},\mathbf{v}_{t}=\mathbf{W}_{V}\mathbf{x}_{t},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\beta_{t}=g(\mathbf{x}_{t})},~~\mathbf{o}_{t}=\sum_{i=1}^{t}\frac{\exp\left({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\beta_{i}^{t-i}}\mathbf{q}_{t}^{\top}\mathbf{k}_{i}\right)}{\sum_{j=1}^{t}\exp\left({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\beta_{j}^{t-j}}\mathbf{q}_{t}^{\top}\mathbf{k}_{j}\right)}\mathbf{v}_{i}.$ (3)
Here, we propose a retention gate $g$, which is a lightweight network, to parametrize the token importance $\beta_{t}$. The retention gate can be a linear projection, i.e., $g(\mathbf{x})=\sigma(\mathbf{W}_{\beta}\mathbf{x}_{t}+b),\mathbf{W}_{\beta}\in\mathbb{R}^{1\times d}$, or a simple MLP, i.e., $g(\mathbf{x})=\sigma(\mathrm{MLP}(\mathbf{x})+b)$. The sigmoid function $\sigma$ squashes the output of $g$ to the range $[0,1]$, while $b$ is a learnable bias. When all $\beta_{t}=1,\forall t$, our retention-gated attention recovers the vanilla attention. Our ablation studies show that an MLP with a single hidden layer provides a more powerful retention estimation than a linear projection.
@@
Brain-inspired Interpretation. Our proposed retention-gated attention bears a natural connection to the classical Ebbinghaus’s forgetting curve theory (ebbinghaus2013image), which models human memory retention as an exponential decay over time. A common approximation of human retention rate is $R=\exp(-tS)$, where $t$ is the time and $S$ is the memory strength determining how fast $R$ decays over time in the absence of training (wozniak1995two).
In a similar spirit, our retention-gated attention models the contribution of past tokens as an exponentially decreasing function of their temporal distance from the current step, i.e., $\exp((t-i)\log\beta_{i})$. Each token begins with full weight ($\mathrm{bar}\alpha_{ii}=1$), akin to a newly encoded memory, and its influence decays as more tokens arrive, mirroring the drop in recall probability described by the forgetting curve. The parameter $\beta$ acts analogously to memory strength $S$: larger values yield more persistent, durable memories, while smaller values indicate weaker memories that fade quickly.
This connection provides an intuitive justification for our design. By embedding a forgetting mechanism into attention, we enable the model to dynamically prioritize recent or intrinsically important tokens while gradually discarding less informative context, mirroring how humans manage limited memory capacity in practice. Note that zhong2024memorybank also drew on Ebbinghaus’s forgetting curve to construct a long-term memory bank, but their focus was on retrieval-augmented generation, whereas our approach integrates forgetting directly into the attention mechanism.

### v2

# arXiv:2512.03324v2[cs.LG] 01 Mar 2026

@@ 1 Introduction
Modern large language models (LLMs) can, in principle, handle extremely long input contexts – some recent models support context windows of 128k tokens or more (Yang et al., 2025; Gao et al., 2024). Yet, extending context length comes with steep computational costs. The self-attention mechanism has quadratic time complexity in sequence length, and storing the key-value (KV) cache for thousands of tokens can quickly exhaust GPU memory (Wang et al., 2025; Li et al., 2024a). In practical deployments, the KV cache, which saves past key and value vectors to avoid re-computation, becomes a major memory and latency bottleneck for long-context inference. Decoupling resource usage from context length is therefore critical for enabling efficient and scalable applications such as long-horizon reasoning (Chen et al., 2025) and lifelong agents (Zheng et al., 2025; Li et al., 2024d).
@@
To address this challenge, recent work has explored memory-bounded LLMs that can operate effectively under constrained KV budgets (Li et al., 2024a). One line of research focuses on compression and quantization, aiming to reduce memory footprint by learning compact representations of past tokens rather than storing all keys and values explicitly (Hooper et al., 2024; Saxena et al., 2024). These techniques are mostly effective during the prefill phase but scale poorly with generation length. Another line leverages attention sparsity to offload most of the cache to CPU or secondary storage, and retrieve only relevant segments on demand via similarity search (Tang et al., 2024) or learned indices (Gao et al., 2025). While offloading lowers the on-GPU footprint, it incurs nontrivial orchestration overhead that accumulates over long generations, undermining end-to-end throughput.
@@
A more common and direct approach to enforce a fixed memory budget is KV cache eviction, which directly drops certain tokens from the KV cache (Xiao et al., 2023). Many KV eviction strategies have been proposed to decide which tokens to remove. However, most of them are attention-guided heuristics: they track attention from new queries to cached tokens and retain those that are recently or frequently attended, adapting the cache to the current focus (Zhang et al., 2023; Li et al., 2024c; Wang et al., 2025; Liu et al., 2025; Ghadia et al., 2025; Cai et al., 2025). While being efficient, these methods assume that recent attention is a reliable proxy for future importance. This assumption often breaks for long-horizon generation and reasoning tasks: a token might be crucial much later, even if it has not been attended to in the recent past (Jiang et al., 2024). Moreover, attention-based eviction can suffer from attention bias, e.g., the model might temporarily overlook a needed token due to a distracting context (Shi et al., 2023), causing it to be evicted prematurely. While some recent studies have attempted to learn better eviction decisions (Chen et al., 2024; Zeng et al., 2024), these methods typically scale poorly with sequence length and are therefore limited to the prefilling stage.
@@
In this work, we take a new perspective on the KV eviction problem. Rather than relying on the attention-guided importance, we propose to learn each token’s intrinsic importance at the time of its creation and use that as the basis for eviction. Intuitively, not all tokens are created equal: some carry significant semantic or task-related weight (e.g. a critical fact, a question being answered, or the first few “sink" tokens that often encode the topic or instructions), while others are relatively inconsequential (e.g. filler words, stopwords, or trivial arithmetic steps in a chain-of-thought). Moreover, the importance of tokens is not uniform across the network, but it varies systematically by layers and heads, reflecting their functional specializations (Voita et al., 2019; Wu et al., 2024b).
@@
We posit that the contextual embedding of a token already encodes much of its long-term utility. We therefore introduce a retention gate that maps the token’s embedding and produces a scalar retention score $\beta\in[0,1]$ reflecting the token’s inherent importance for a specific layer and head. Especially, we design this retention score to decay exponentially as the context grows, mimicking the gradual forgetting of old information in human brains (Ebbinghaus, 2013). Thus, a highly important token will have $\beta\approx 1$ and retain a high score for a long time, whereas a token deemed unimportant will have $\beta$ closer to 0 and its influence will vanish quickly. We leverage this score to drive a simple eviction policy: whenever the number of cached tokens exceeds the budget $M$, we evict the token with the smallest current retention score. This approach, which we call Token RetentIon for Memory-bounded KV Cache (TRIM-KV), ensures that at all times, the cache is filled with the $M$ tokens judged most intrinsically important, with a preference toward more recently generated tokens.
@@
Results and Contributions. Through extensive experiments on long-context and long-generation benchmarks, we demonstrate that our learnable token retention approach substantially improves the performance of memory-bounded LLMs. On challenging mathematical reasoning datasets, GSM8K, MATH, AIME, a long procedural generation benchmark, LongProc, and a long-memory chat assistant benchmark, LongMemEval, our method consistently outperforms eviction baselines, even when those baselines use $4\times$ more KV budget, and deliver 58.9% relative gain on pass@1 compared to the SOTA learnable KV retrieval baseline (Gao et al., 2025), especially in low-memory regimes. Remarkably, in several settings, TRIM-KV even surpasses a full-cache model, suggesting that selective retention can function as an effective regularizer by suppressing noise from uninformative tokens.
We also present qualitative evidence that learned retention scores align with human intuition: the model tends to assign high scores to initial tokens and problem descriptions, and low scores to less meaningful punctuation. Notably, many behaviors reminiscent of common heuristics, such as keeping sink tokens, sliding windows, and gist compression (Mu et al., 2023), emerge naturally and adaptively from our learned policy, without being hard-coded. Finally, we show that these learned retention scores can also act as a diagnostic tool for probing layer- and head-specific dynamics, providing a lightweight means to analyze and ultimately improve the interpretability of attention patterns.
@@ 2 Related Work
KV Cache Compression. As model sizes and context windows grow, optimizing KV-cache memory is increasingly critical. Prior work largely falls into three directions: (i) token eviction/merging (Xiao et al., 2023; Li et al., 2024c; Zhang et al., 2023; Nawrot et al., 2024; Zhang et al., 2024; Qin et al., 2025; Wang et al., 2025; Liu et al., 2025; Park et al., 2025; Cai et al., 2025; Park et al., 2025; Kim et al., 2024), (ii) vector compression/quantization (Hooper et al., 2024; Liu et al., 2024b; Yue et al., 2024; Sun et al., 2024a), and (iii) token retrieval (Tang et al., 2024; Liu et al., 2024a; Gao et al., 2025). While effective in many settings, vector compression and retrieval either discard fine-grained information or introduce nontrivial systems overhead (e.g., coordination and data movement) (Li et al., 2024a). Moreover, their memory and computation still scale with sequence length, making them inefficient for long-horizon generation applications. Token eviction offers a simple, memory-bounded alternative; however, most existing policies are heuristic and can significantly degrade performance, especially on long reasoning trajectories. Recent work has introduced learnable eviction policies (Chen et al., 2024; Zeng et al., 2024; Huang et al., 2024), but these are primarily designed for the pre-filling stage and thus are not well suited to sustained long-horizon generation. We bridge this gap by introducing a learnable and efficient eviction policy designed for long-horizon LLM inference under fixed memory budgets.
Forgetting in Language Models. A key limitation of vanilla self-attention is the lack of an explicit forgetting mechanism, forcing the model to carry potentially irrelevant information and making long-context processing inefficient. Early work tackled this by replacing quadratic attention with linearized and recurrent variants (Katharopoulos et al., 2020; Wang et al., 2020; Sun et al., 2023; Yang et al., 2023; 2024) that summarize the past into a fixed-size state, often a single vector. While computationally attractive, such heavy compression can degrade performance on tasks requiring long-range memory. Follow-up studies (Behrouz et al., 2024; Sun et al., 2024b; Karami et al., ; Karami and Mirrokni, 2025) increase memory capacity by replacing this hidden vector with a more expressive neural state. Complementary lines of work retain softmax attention but enforce forgetting by modifying attention logits (Lin et al., 2025) or imposing trainable sparsity patterns (Yuan et al., 2025). However, these approaches typically alter attention dynamics substantially and thus require training models from scratch. This incurs significant training cost and leaves their scalability to contemporary LLM sizes uncertain. In contrast, we introduce a plug-in forgetting mechanism for pretrained LLMs that converts them into memory-bounded models, providing long-context efficiency without retraining from scratch.
@@ 3.1 Transformers with Self-Attention
where $\mathbf{q},\mathbf{k},\mathbf{v}$ are query, key, and value states, respectively, and $\mathbf{W}_{Q},\mathbf{W}_{K},\mathbf{W}_{V}\in\mathbb{R}^{d\times d}$ are linear transformation weights. Here, we assume a single-head attention layer and omit the scaling factor $1/\sqrt{d}$ for simplicity. The sequence of key-value pairs $\{(\mathbf{k}_{i},\mathbf{v}_{i})\}_{i}$ is the in-context memory of the LLM. During the autoregressive decoding, we typically generate one token at a time and cache the running key-value pair $(\mathbf{k}_{t},\mathbf{v}_{t})$ to our in-context memory to avoid recomputation. However, this vanilla caching approach leads to a linear increase in memory footprint with the sequence length, while computation grows quadratically (Keles et al., 2023). This reduces efficiency when handling long-context inputs and extended generation tasks.
@@ 3.2 Revisiting KV Cache Eviction
$\mathbf{o}_{t}^{\prime}=\sum_{i=1}^{t}\frac{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\alpha_{ti}}\exp\left(\mathbf{q}_{t}^{\top}\mathbf{k}_{i}\right)}{\sum_{j=1}^{t}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\alpha_{tj}}\exp\left(\mathbf{q}_{t}^{\top}\mathbf{k}_{j}\right)}\mathbf{v}_{i}\quad\text{where}\quad~\alpha_{ti}\in\{0,1\}~\text{and}~\alpha_{ti}\geq\alpha_{t+1,i},~~\forall i,t.\vskip-11.38109pt$ (1)
@@
Solving the above constrained optimization at every time step $t$ is impractical due to its combinatorial nature and efficiency requirements of LLM inference in real-world applications. Most existing approaches  (Xiao et al., 2023; Han et al., 2023; Zhang et al., 2023; Li et al., 2024c; Cai et al., 2025; Ghadia et al., 2025) opt to determine $\alpha$ heuristically while we focus on a learnable eviction method.
@@ 4.1 Selective In-Context Memory via Retention-Gated Attention
We introduce retention-gated attention, a trainable mechanism that mimics the information loss induced by inference-time eviction. From the formulation (1), the sequence $\alpha_{ii},\alpha_{(i+1)i},\ldots,\alpha_{ti}$ represents how token $i$ is retained in the attention computation over time. Retention begins at 11 and then abruptly drops to 0 once the token is evicted. While this binary behavior matches the inference stage, it poses challenges for learning: the signal is discrete, non-differentiable, thus providing no gradients for optimization. To remedy this, we replace the hard binary variable $\alpha$ with a smooth, monotonically decreasing function that models the gradual decay of importance while enabling gradient-based training. A natural candidate is the sigmoid function, $\mathrm{bar}{\alpha}_{ti}=1/(1+\exp(f(\mathbf{x}_{i},t)))$, which predicts the time at which the token is evicted. However, this design suffers from two drawbacks: (i) the domain of $f$ is unnormalized since the sequence length is unknown during decoding, and (ii) the sigmoid flattens across most of its range, producing negligible variation between steps and leading to vanishing gradients during training.
@@
$\mathbf{q}_{t}=\mathbf{W}_{Q}\mathbf{x}_{t},\mathbf{k}_{t}=\mathbf{W}_{K}\mathbf{x}_{t},\mathbf{v}_{t}=\mathbf{W}_{V}\mathbf{x}_{t},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\beta_{t}=g(\mathbf{x}_{t})},~~\mathbf{o}_{t}=\sum_{i=1}^{t}\frac{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\beta_{i}^{t-i}}\exp\left(\mathbf{q}_{t}^{\top}\mathbf{k}_{i}\right)}{\sum_{j=1}^{t}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\beta_{j}^{t-j}}\exp\left(\mathbf{q}_{t}^{\top}\mathbf{k}_{j}\right)}\mathbf{v}_{i}.$ (3)
Here, we propose a retention gate $g$, which is a lightweight network, to parametrize the token importance $\beta_{t}$. The retention gate can be a linear projection, i.e., $g(\mathbf{x})=\sigma(\mathbf{W}_{\beta}\mathbf{x}_{t}+b),\mathbf{W}_{\beta}\in\mathbb{R}^{1\times d}$, or a simple MLP, i.e., $g(\mathbf{x})=\sigma(\mathrm{MLP}(\mathbf{x})+b)$. The sigmoid function $\sigma$ squashes the output of $g$ to the range $[0,1]$, while $b$ is a learnable bias. When all $\beta_{t}=1,\forall t$, our retention-gated attention recovers the vanilla attention. By incorporating the retention score into the exponential term, the attention weight can be written as $\exp(\mathbf{q}_{t}^{\top}\mathbf{k}_{i}+(t-i)\log\beta_{i})$, which reveals that the retention score acts as an additive bias on the attention logits (Press et al., 2021).
@@
Brain-inspired Interpretation. Our retention-gated attention bears a natural connection to the Ebbinghaus’s forgetting curve (Ebbinghaus, 2013), which models human memory retention as exponential decay, $R=\exp(-tS)$, where $S$ denotes memory strength (Woźniak et al., 1995).
Similarly, our retention-gated attention models each token’s contribution as exponentially decaying with temporal distance, $\exp((t-i)\log\beta_{i})$. Tokens start with full weight and gradually lose influence as new tokens arrive. The parameter $\beta$ acts analogously to memory strength: larger values yield more persistent, durable memories, while smaller values indicate weaker memories that fade quickly. This design embeds a forgetting mechanism into attention, enabling dynamic prioritization of recent or salient tokens while discarding less useful context, akin to human memory. Unlike prior work that applies the forgetting curve to retrieval-augmented generation (Zhong et al., 2024), we integrate it directly into the attention mechanism.